OpenStudy (anonymous):

*Fan & medal* Use the quadratic model P(x) = 25x^2 - 24x + 615 to predict P(x) if x equals 8. A. 2203 B. 1672 C. 2023 D. 2677

OpenStudy (anonymous):

Please help, I don't know how to do this. ;c

OpenStudy (mathmale):

All you have to do is to let x = 8 in the function P(x) = 25x^2 - 24x + 615.

OpenStudy (jordanloveangel):

Simplifying P(x) = 25x2 + -24x + 615 Multiply P * x xP = 25x2 + -24x + 615 Reorder the terms: xP = 615 + -24x + 25x2 Solving xP = 615 + -24x + 25x2 Solving for variable 'x'. Reorder the terms: -615 + 24x + xP + -25x2 = 615 + -24x + 25x2 + -615 + 24x + -25x2 Reorder the terms: -615 + 24x + xP + -25x2 = 615 + -615 + -24x + 24x + 25x2 + -25x2 Combine like terms: 615 + -615 = 0 -615 + 24x + xP + -25x2 = 0 + -24x + 24x + 25x2 + -25x2 -615 + 24x + xP + -25x2 = -24x + 24x + 25x2 + -25x2 Combine like terms: -24x + 24x = 0 -615 + 24x + xP + -25x2 = 0 + 25x2 + -25x2 -615 + 24x + xP + -25x2 = 25x2 + -25x2 Combine like terms: 25x2 + -25x2 = 0 -615 + 24x + xP + -25x2 = 0 The solution to this equation could not be determined. i think this is how to solve it

OpenStudy (jhannybean):

What do you mean by `xP`?

OpenStudy (anonymous):

Ok so I just do this? P(x) = 25*8^2 - 24*8 + 615???

OpenStudy (anonymous):

If so the answer would be C.2023 right??? I think I did that right, I'm not that good with these kinda problems. :/

OpenStudy (mathmale):

Yes. The formula P(x) = 25x^2 - 24x + 615 enables you to predict P(x) for any x in its domain. This is a polynomial function, so you may substitute 8 for x (8 is in the domain). The result is a prediction / estimate for the value of P(x) when x= 8.

OpenStudy (anonymous):

Thanks, this helps a ton. I was getting confused on this question.